A particular enzyme at a research facility is being studied by a group of graduate students. This enzyme has a Km value of 5.0 X 10-6 M. The students study this enzyme with an initial substrate concentration of 0.055 M. At one minute, 7 μM of product was made. What is the amount of product produced after 5 minutes?What is the Vmax?(Hint: unit of velocity is μM/min)


1
Expert's answer
2021-02-01T11:21:15-0500

Vo=Vmax[Substrate]/{Km+[Substrate]}

7.0*10-6M=Vmax(0.055M)/{5.0*10-6M +0.055M}

Vmax=7.1*10-6M/min

At 5 minutes, the amount of product formed is:

[P]=Vmax*t

7.1*10-6M/min*5min

[P]=3.55*10-5M